OpenStudy (lukecrayonz):

Equation of the tangent line of the graph 8x/(x^2+1) at the origin and (-1,-4) These are two separate answers

OpenStudy (lukecrayonz):

Nevermind I know how :)

OpenStudy (lukecrayonz):

@kc_kennylau actually, I got it for (-1,-4) but not for the origin!

OpenStudy (lukecrayonz):

nevermind, I think it's 8x and 4x!

OpenStudy (kc_kennylau):

sorry for correcting, but never mind is two words

OpenStudy (lukecrayonz):

Haha I honestly did not know that..

OpenStudy (lukecrayonz):

If:\[\log_{b}4=1.386\] \[\log_{b}7=1.946\] what is\[\log_{b}28\]

OpenStudy (lukecrayonz):

Just 1.386*1.946?

OpenStudy (kc_kennylau):

Yep

OpenStudy (lukecrayonz):

How do I find velocity and acceleration when only given speed?

OpenStudy (kc_kennylau):

Maybe not related, but this is interesting: Unit of distance: \(\large m\cdot t^0\) Unit of speed (velocity): \(\large m\cdot t^{-1}\) Unit of acceleration: \(\large m\cdot t^{-2}\)

OpenStudy (kc_kennylau):

but speed is the same as velocity?

OpenStudy (lukecrayonz):

Idk ;_;

OpenStudy (lukecrayonz):

http://gyazo.com/79d312d9c3243727e7c78786a5130f74

OpenStudy (lukecrayonz):

@Kainui

OpenStudy (kc_kennylau):

s is not speed, s is displacement (distant)

OpenStudy (lukecrayonz):

Clearly I'm dumb and it's 4 a.m :P

OpenStudy (kc_kennylau):

Important formulas: \[\frac{ds}{dt}=v\]\[\frac{dv}{dt}=a\]

OpenStudy (kc_kennylau):

So I assume you know how to do that? :P

OpenStudy (lukecrayonz):

So basically a is second derivative

OpenStudy (lukecrayonz):

That's what I figured it was but I didn't want to mix them up

OpenStudy (kc_kennylau):

yep :)

OpenStudy (lukecrayonz):

How do you find the derivative of (lnx)^2?

OpenStudy (lukecrayonz):

2lnx? that's it? haha i mean im sure theres more steps

OpenStudy (kc_kennylau):

Chain rule :)

OpenStudy (lukecrayonz):

2 * ln x * (1/x)

OpenStudy (lukecrayonz):

I hate quotient rule xD I actually hate all rules besides power

OpenStudy (kc_kennylau):

sorry back

OpenStudy (kc_kennylau):

yep

OpenStudy (kc_kennylau):

but you have to simplify it

OpenStudy (lukecrayonz):

2lnx/x

OpenStudy (kc_kennylau):

sorry back yep

OpenStudy (lukecrayonz):

And then 2ln(8)/8=0.519

OpenStudy (kc_kennylau):

yep

OpenStudy (lukecrayonz):

http://gyazo.com/c45644d8437428e85dfd85b9e056275d

OpenStudy (lukecrayonz):

So then input y=0.519...x+b and input the (8,0) for x and y, solve for b and thats my y intercept

OpenStudy (lukecrayonz):

Just not sure how i'd do this without a computer since the numbers are extremely long

OpenStudy (kc_kennylau):

you have a calculator

OpenStudy (kc_kennylau):

you don't need so many places. 4 is actually enough.

OpenStudy (lukecrayonz):

Would you say the 2nd part of B is asking for the derivative of P'(x) when x=70 as in, no delta included, or with the delta? http://gyazo.com/798ed926afb065b09f95a5675a583017

OpenStudy (lukecrayonz):

@kc_kennylau

OpenStudy (kc_kennylau):

\[\Delta P\approx P'(x)\cdot\Delta x\]

OpenStudy (lukecrayonz):

Wait, I did it wrong -_- forgot about R(x)

OpenStudy (kc_kennylau):

i see

OpenStudy (lukecrayonz):

So the derivative of P'(x) is -0.02x+1.4

OpenStudy (lukecrayonz):

Of P(x) not of P'(x)

OpenStudy (kc_kennylau):

sorry back

OpenStudy (lukecrayonz):

I suck at these with delta

OpenStudy (kc_kennylau):

what's P(x) anyway

OpenStudy (kc_kennylau):

brb

OpenStudy (lukecrayonz):

-0.01x^2+1.4x-50

OpenStudy (kc_kennylau):

yep

OpenStudy (lukecrayonz):

So to find delta P, I do what?

OpenStudy (kc_kennylau):

\[\Delta P\approx P'(x)\cdot\Delta X\]

OpenStudy (kc_kennylau):

sorry brb

OpenStudy (lukecrayonz):

-0.02x+1.4*1..?

OpenStudy (lukecrayonz):

-0.02(70)+1.4*1..?

OpenStudy (kc_kennylau):

brackets bro, brackets

OpenStudy (kc_kennylau):

[-0.02(70)+1.4]*1

OpenStudy (kc_kennylau):

well anyway theyre equal

OpenStudy (kc_kennylau):

sorry really be right back

OpenStudy (lukecrayonz):

Haha it's okay im getting off, 5:30 a.m it's time for me to take a nap then get up and study some more for my final!

OpenStudy (kc_kennylau):

hi lol back